# ContactHandover: Contact-Guided Robot-to-Human Object Handover

### v1

@@ V RESULTS AND ANALYSIS
ContactHandover achieves the most successful handovers. As shown in Table I, ContactHandover achieves visible and reachable handovers for all objects in all runs, with an average success rate of 68.5%. From our ablations, estimating the handover position (Ablation 3) contributes the most to improvements in overall success rate (50% improvement), and in particular, the reachability (70% improvement), compared to no optimization (Ablation 4). Ablation 4 does not achieve any successful handovers as all of the objects fall out of reach range of the human arm.
Our contact-guided grasp selection and handover orientation estimation algorithms further improves handover visibility and reachability. As a result, ContactHandover improves the final success rate by 18.5% compared to Ablation 3.
Grasp re-ranking yields more available human contacts during handover. We show the effect of contact-guided grasp selection by comparing with Ablation 1, shown in Fig. 3(b). Without considering human contacts when selecting stable grasps, the robot gripper will often block human contact points on the object. As shown in Fig 3(b), the robot grasps objects like hammer and pan by the handles. While it orients the human contact points towards the human during delivery, the robot’s gripper blocks human from receiving the objects on the handles. Therefore the human contacts are both less visible due to occlusion from gripper and less reachable due to less contact points between gripper and human.
We note that there are cases where grasp re-ranking makes little effect to the final handover. For example, if all stable robot grasps overlap with human contacts, for instance, on the handle of the knife, the selected robot grasp will block a large portion of human contacts regardless. On the other hand, if the majority of stable robot grasps do not overlap with the predicted human contacts, penalizing human contact occlusion makes little difference. For instance, most stable robot grasps on a mug are on its rim, rather than the handle where human prefers to grasp. Nonetheless, qualitatively, on all objects, ContactHandover on average outperforms Ablation 1 by 5.5%.
Estimating handover orientation makes more human contacts face the receiver. We show the effect of handover orientation estimation by comparing with Ablation 2, shown in Figure 3(c). In Ablation 2, the robot does not orient contact points towards the human during delivery. Although the robot gripper leaves out the human contact parts while grasping, human contact parts are still inaccessible to the human in the final handover. For instance, human-preferred contacts could be on the opposite side of the robot gripper from the human, making them unreachable (e.g. scissors, hammer and knife handles); they could also be self-occluded from the human’s point of view (e.g. hammer, mug and pan handles). Overall, ContactHandover on average outperforms Ablation 2 by 17.4%.
@@
Limitations. Our work leverages a hand-object contact dataset to learn a proxy for human receiving preferences. However, there are a few limitations to this approach. Firstly, our human contact predictions are inferred only from object shape. In certain scenarios, the object state or function may also matter, and cannot be simply inferred from its shape. For example, if a container such as mug or bowl has water inside, the object should be kept its canonical pose (opening facing up) during the grasping and delivery phase to prevent spilling. Secondly, our human contact dataset is limited in scale. While the dataset contains 27 diverse household objects, it’s still a small fraction of objects we can expect a human or robot to interact with in daily life. Future work could consider expanding the dataset by collecting hand-object contacts for more objects. We believe learning from a larger human contact dataset can help increase our method’s generalization ability to unseen objects.
@@ VI CONCLUSIONS
We propose ContactHandover, a two-phase robot to human handover system guided by human contacts. The robot first achieves a stable grasps the object while accommodating human contact preferences through a grasp pose re-ranking mechanism. Then the robot delivers the object to a pose that minimizes human arm joint torques and displacements and maximizes available human preferred contacts to the receiver. To evaluate our system, we propose two quantitative metrics that measure the visibility and reachability of an object’s human preferred contacts during handover. We evaluate our system on 27 diverse household objects and demonstrate that our system achieves better visibility and reachability of the predicted human contact areas to the receiver comparing to several ablations.

### v2

@@ V RESULTS AND ANALYSIS
ContactHandover achieves the most successful handovers. As shown in Table I, ContactHandover achieves visible and reachable handovers for all objects in all runs, with an average success rate of 68.5%. From our ablations, estimating the handover position (A3) contributes the most to improvements in overall success rate (0% to 50%), and in particular, the reachability (0% to 70%), compared to no optimization (A4). Ablation 4 does not achieve any successful handovers as all of the objects fall out of reach range of the human arm.
Our contact-guided grasp selection and handover orientation estimation algorithms further improves handover visibility and reachability. ContactHandover improves the final success rate by 18.5 percentage points compared to A3.
Grasp re-ranking yields more available human contacts during handover. We show the effect of contact-guided grasp selection by comparing with Ablation 1, shown in Fig. 3(b). Without considering human contacts when selecting stable grasps, the robot gripper will often block human contact points on the object. As shown in Fig 3(b), the robot grasps objects like hammer and pan by the handles. While it orients the human contact points towards the human during delivery, the robot’s gripper blocks human from receiving the objects on the handles. Therefore, the human contacts are less visible due to occlusion from gripper and less reachable due to less contact points between gripper and human.
We note that there are cases where grasp re-ranking makes little effect to the final handover. For example, if all stable robot grasps overlap with human contacts, for instance, on the handle of the knife, the selected robot grasp will block a large portion of human contacts regardless. On the other hand, if the majority of stable robot grasps do not overlap with the predicted human contacts, penalizing human contact occlusion makes little difference. For instance, most stable robot grasps on a mug are on its rim, rather than the handle where human prefers to grasp. Nonetheless, qualitatively, on all objects, Grasp re-ranking increases success rate from 63% (A1) to 68.5% (ContactHandover).
Estimating handover orientation faces more human contacts to the receiver. We show the effects of handover orientation estimation by comparing with Ablation 2, shown in Fig. 3(c). In Ablation 2, the robot does not orient contact points towards the human during delivery. Although the robot gripper leaves out the human contact parts while grasping, human contact parts are still inaccessible to the human in the final handover. For instance, human-preferred contacts could be on the opposite side of the robot gripper from the human, making them unreachable (e.g. scissors, hammer and knife handles); they could also be self-occluded from the human’s point of view (e.g. hammer, mug and pan handles). Estimating handover orientation, on average, increases success rate from 51.1% (A2) to 68.5% (ContactHandover).
@@
Limitations. Our work leverages a hand-object contact dataset to learn a proxy for human receiving preferences. However, there are a few limitations to this approach. Firstly, our human contact predictions are inferred only from object shape. In certain scenarios, the object state or function may also matter, and cannot be simply inferred from its shape. For example, if a container such as mug or bowl has water inside, the object should be kept its canonical pose (opening facing up) during the grasping and delivery phase to prevent spilling. Secondly, our human contact dataset is limited in scale. While the dataset contains 27 diverse household objects, it’s still a small fraction of objects we can expect a human or robot to interact with in daily life. Future work could consider expanding the dataset by collecting hand-object contacts for more objects or leveraging multimodal foundation models [12, 24, 25]. We believe learning from a larger human contact dataset can help increase our method’s generalization ability to unseen objects. Lastly, in our evaluation, we assume a standard human pose in ‘standing’ mode, and that the human receives the object only after the robot delivers the object to the desired pose. Therefore, the policy is not reactive to real-time human pose changes and movements. Future work could consider combining ContactHandover with human pose estimation and hand motion tracking.
@@ VI CONCLUSIONS
We propose ContactHandover, a two-phase robot to human handover system guided by human contacts. The robot first achieves a stable grasp on the object while accommodating human contact preferences through a grasp pose re-ranking mechanism. Then the robot delivers the object to a pose that minimizes the human arm joint torques and displacements, as well as the total distances from contact points to the receiver. To evaluate our system, we propose two quantitative metrics that measure the visibility and reachability of an object’s human preferred contacts during handover. We evaluate our system on 27 diverse household objects, demonstrating improved visibility and reachability of predicted human contact areas compared to several ablations.
ACKNOWLEDGMENT
This work was supported in part by the NSF Award #2037101, and #2132519. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the sponsors.